Sketch the level curves f
1(c) for the following functions:
f(x, y, z) = x-y2-z2
, c = 1, 0, 1
1
Expert's answer
2020-10-19T17:20:10-0400

We have been provided the map


"f(x,y,z)=x-y^2-z^2"

Now, level curve is defined as

"G_f=\{(x,y,z)\in \mathbb{R}^3|f(x,y,z)=c\}"

That is, here, "f^{-1}(c)=G_f"

Now, set

"f(x,y,z)=c\implies (x-c)=y^2+z^2"

Clearly, the level curves are family of paraboloid which passes through the center "(c,0,0)"











The violet, green and red are the sketch of level curve of "f^{-1}(-1),f^{-1}(0)\& f^{-1}(1)" respectively.